OpenStudy (anonymous):

a) What is the slope of the line through the points (5, 2) and (5, -3)? b) Describe the line through these points.

OpenStudy (anonymous):

a) Change in y's over change in x's

OpenStudy (anonymous):

\[\frac{ y2-y1 }{ x2-x1 }\] \[\frac{ -3-2 }{ 5-5 } = \frac{ -5 }{ 0 }\]

OpenStudy (anonymous):

b) Its a verticle line